OpenStudy (anonymous):

1. A farmer has cows and ducks on a truck. These are the only creatures on the truck. Altogether, there are 15 heads and 44 legs on the truck. How many cows are on the truck? How many ducks?

OpenStudy (anonymous):

x is the cows y is the ducks x+y = 15 Cows have 4 legs ducks have 2 4x+2y=44 Multiply the first equation by 2 2x+2y =30 4x+2y=44 -2x=-14 Divide each side by -2 x=7 Plug x=7 into the first equation, x+y=15 7+y=15 y=8 So there are ? cows and ? ducks. Can you answer that question? The "So there are ____ cows and ____ ducks." part? :)

OpenStudy (anonymous):

Trying to figure this out give me a sec :)

OpenStudy (anonymous):

Okay :).

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

Ok I kinda understand what our doing I just still dont knowhow to figure its just a little hard to understand what you put

OpenStudy (anonymous):

Well x = cows and y = ducks. I told you the varriable y and x equal something. That something was x = 7 y = 8 Equation: 7(cows) + 8(ducks) = 15